# Thermalization and Many-Body Zeno Effect in monitored Hamiltonian Dynamics

### v1

@@ V Zeno effect
Excessively frequent measurements can freeze the quantum state due to the Zeno effect, thereby suppressing its inherent randomness. For a finite evolution time $T$, if the number of measurements $n$ is too large, the bath will remain almost fixed in its reset state $|\phi_{0}\rangle$. Consequently, the probability of obtaining any measurement outcome other than the reset state approaches zero. The system therefore remains in a pure state, with the frame potential approaching one.
@@
$F^{(K)}\gtrsim e^{-O\left(\frac{T^{\alpha+1}}{n^{\alpha}}\right)},$ (12)
where $\alpha=1$ is a constant.
[…]
When $n$ is large, the frame potential approaches 1. To extract the exponential decay behavior, we plot the data on a logarithmic scale. The asymptotic behavior of the frame potential for different Hamiltonians is shown in Fig. 5. The numerical results are represented by the blue curves, while the exponential decay rate given in Eq. (12) is indicated by the green lines. As $n$ increases, the numerical results converge more closely to the predicted exponential behavior. In the case where $\left[H_{{\rm c}0},|\phi_{m}\rangle\langle\phi_{m}|\right]=0$, the decay exponent is $\alpha=3$, as illustrated in Fig. 5(a). Otherwise, when the commutator is nonzero, the exponent is $\alpha=1$, as shown in Fig. 5(b) and (c).
The starting point for the number of measurements required to observe the Zeno effect is approximately
@@
where $c_{H}$ is a constant arising from the Hamiltonian, and its value will be derived in Appendix B. The number of measurements should not exceed this value to preserve states with a good level of randomness.
The saturation of the frame potential requires that $n_{\rm sat}<n_{\rm Zeno}$. In other words, the total evolution time must satisfy
$T>\sqrt[\alpha+1]{\frac{(n_{\rm s}-{\rm log}_{2}r)^{\alpha}n_{\rm s}\ln 2}{n_{\rm b}^{\alpha}c_{H}}}\leavevmode\nobreak\ .$ (14)
@@
When the total evolution time $T$ is sufficiently short such that $n_{\rm sat}>n_{\rm Zeno}$, there is a competition between HDT and Zeno effect. In this regime, the frame potential does not reach its saturation value and remains significantly above it. The minimum frame potential is attained at around
@@
Figure 5: Frame potential as a function of the number of measurements for (a) $H_{\rm Ising}$, (b) $H_{YY}$, and (c) $H_{XXX}$. Blue lines represent numerical results, while green lines correspond to analytical predictions for Zeno effect with (a) $\alpha=3$, and (b)(c) $\alpha=1$. The total evolution time is $T=15$. The system consists of 5 qubits, with 3 bath qubit.
VI conclusion
We employ a protocol involving multiple Hamiltonian evolutions, measurements, and resets to access holographic deep thermalization. To quantify the degree of randomness generated, we use the frame potential. In the regime of few measurements, the frame potential decreases exponentially until it saturates at a minimum value, indicating maximal scrambling. In contrast, when measurements are too frequent, the frame potential rises exponentially because the quantum Zeno effect suppresses thermalization. We derive the asymptotic behavior of the frame potential analytically, which shows excellent agreement with our numerical results.
@@ Acknowledgements.
The authors acknowledges discussions with Bingzhi Zhang. The project is supported by Office of Naval Research Grant No. N00014-23-1-2296 and DARPA HR0011-24-9-0362. QZ also acknowledges support from NSF (OMA-2326746, 2350153, CCF-2240641), AFOSR MURI FA9550-24-1-0349, DARPA HR0011-24-9-0362, Halliburton Company an unrestricted gift from Google.
@@ Appendix B Zeno effect
In the following, we derive Eq. (12). The Hamiltonian can be decomposed into three parts
@@
$=H_{{\rm s}0}+H_{{\rm b}}+H_{{\rm c}0},$ (18)
where $H_{{\rm s}},H_{{\rm b}}$ and $H_{{\rm c}}$ corresponds to the system, bath and interaction Hamiltonian. In the second step, we introduced a modified system and interaction Hamiltonian
@@
$H_{{\rm c}0}=H_{\rm c}-\langle\phi_{0}|H_{\rm c}|\phi_{0}\rangle\leavevmode\nobreak\ ,$ (20)
@@ B.1 Low order Zeno effect
In the general case, the powers is $\alpha=1$. The measurement basis are changed under the operation $H_{{\rm c}0}$, Namely $\left[H_{{\rm c}0},|\phi_{m}\rangle\langle\phi_{m}|\right]\neq 0$. We can decompose the system wave function in the orthogonal basis
@@
The probability of the bath remain in state $|\phi_{0}\rangle$ is
$P_{0}=1-\sum_{m\neq 0}P_{m}\approx 1-\Delta t^{2}\sum_{l^{\prime}}\overline{|\omega_{l^{\prime},m}|^{2}}.$ (24)
@@
$>$ ${\rm Exp}\left(-\frac{2T^{2}}{n}\sum_{l^{\prime}}\overline{|\omega_{l^{\prime},m}|^{2}}\right).$
In this case, the coefficient is $c_{H}=2\sum_{l^{\prime}}\overline{|\omega_{l^{\prime},m}|^{2}}$.
B.2 Hight order Zeno effect
In the special case, the powers is $\alpha=3$. The measurement basis $|\phi_{m}\rangle$} is invariant under the action of $H_{\rm c0}$—that is, when $\left[H_{\rm c0},|\phi_{0}\rangle\langle\phi_{0}|\right]=0$—we can derive a tighter lower bound on the frame potential.
To facilitate this derivation, we also decompose the bath Hamiltonian into a component that is diagonal in the measurement basis and a remainder term.
$H_{{\rm b}}=H_{{\rm bm}}+V,$ (26)
@@ B.2 Hight order Zeno effect
In the limit of short time evolution $\Delta t\rightarrow 0$, the unitary could be written as [18]
@@
$|\Psi\rangle$ $\approx$ $e^{-iH_{{\rm c0}}\frac{\Delta t}{2}}e^{-iH_{{\rm s0}}\Delta t-iH_{{\rm b}}\Delta t}e^{-iH_{{\rm c0}}\frac{\Delta t}{2}}|\psi_{t}\rangle|\phi_{0}\rangle$ (28)
@@
$|\Psi\rangle\approx\sum_{m}e^{-i\langle\phi_{m}|H_{\rm c0}|\phi_{m}\rangle\frac{\Delta t}{2}}c_{m}|\psi_{t+\Delta t}\rangle|\phi_{m}\rangle.$ (29)
In the first-order time-dependent perturbation theory, we consider only the states directly coupled to the initial state. The corresponding amplitude is given by
@@
$P_{m}$ $\approx$ $4\left|\frac{\langle\phi_{m}|V|\phi_{0}\rangle}{E_{0}-E_{m}}\right|^{2}\sin^{2}\frac{\left(E_{0}-E_{m}\right)\Delta t}{2}.$ (30)
@@
where $|\psi_{T}\rangle$ is the system wave function, and the measurement results remain $|\phi_{0}\rangle$ throughout the entire process.
We can define the corrected factor
$P_{m}^{\prime}>P_{m}\cos^{2K}\frac{\mu_{m}\Delta t}{2},$ (32)
@@
where $2\mu_{m}$ is the difference between the largest and smallest eigenvalues of $\langle\phi_{m}|H_{\rm c0}|\phi_{m}\rangle$.
The frame potential
$F^{(K)}$ $>$ $\left(\sum_{m}P_{m}^{\prime}\right)^{2n}$ (33)
$>$ $\left(1-\frac{1}{4}\sum_{m}\left|\langle\phi_{m}|V|\phi_{0}\rangle\right|^{2}K\mu_{m}^{2}\Delta t^{4}\right)^{2n}$
$>$ ${\rm Exp}\left(-\frac{KT^{4}}{2n^{3}}\sum_{m}\left|\langle\phi_{m}|V|\phi_{0}\rangle\right|^{2}\mu_{m}^{2}\right).$
In this case, the coefficient is $c_{H}=(K/2)\sum_{m}\left|\langle\phi_{m}|V|\phi_{0}\rangle\right|^{2}\mu_{m}^{2}$. Notice that $\mu_{m}=0$ if $\langle\phi_{m}|H_{\rm c0}|\phi_{m}\rangle=0$. Consequently, only a few terms contribute to the summation.

### v2

@@ V Zeno effect
The coupling of the system to the bath can give rise to the ZE [Maimbourg2021, Heinrich2020]. When measurements on the bath are performed excessively often, the quantum ZE can effectively freeze the bath, suppressing its intrinsic randomness. For a finite evolution time $T$, if the number of measurements $n$ is too large, the bath will remain almost fixed in its reset state $|\phi_{0}\rangle$. Consequently, the probability of obtaining any measurement outcome other than the reset state approaches zero. The system therefore remains in a pure state, with the frame potential approaching one.
@@
$F^{(K)}\gtrsim e^{-O\left(\frac{T^{\alpha+1}}{n^{\alpha}}\right)}\penalty 10000\ ,$ (9)
where $\alpha$ is a constant.
In general, the constant $\alpha=1$; while a tighter bound with $\alpha=3$ arises in the presence of a special symmetry. The corresponding deviations are provided in the Appendix B. When $n$ is large, the frame potential approaches 1. To extract the exponential decay behavior, we plot the data on a logarithmic scale. The asymptotic behavior of the frame potential for different Hamiltonians is shown in Fig. 5. The numerical results are represented by the blue curves, while the exponential decay rate given in Eq. (9) is indicated by the green lines. As $n$ increases, the numerical results converge more closely to the predicted exponential behavior.
Combining Eq. (6) with Eq. (9), we can estimate the measurement threshold at which the ZE becomes observable. This threshold is approximately
@@
where $c_{H}$ is a characteristic constant of the Hamiltonian associated with the wave function’s diffusion and scrambling rates. It quantifies the ability of the system’s back-action on the bath. The derivation of $c_{H}$ is provided in the Appendix B. The number of measurements should not exceed this value to preserve states with a good level of randomness.
Saturation of the frame potential requires that it occur before the onset of the ZE, i.e., $n_{\rm sat}<n_{\rm Zeno}$. In other words, the total evolution time must satisfy
$T>\sqrt[\alpha+1]{\frac{(n_{\rm s}-{\rm log}_{2}r)^{\alpha}n_{\rm s}\ln 2}{n_{\rm b}^{\alpha}c_{H}}}\penalty 10000\ .$ (11)
@@
When the total evolution time $T$ is sufficiently short such that $n_{\rm sat}>n_{\rm Zeno}$, there is a competition between HDT and ZE. In this regime, the frame potential does not reach its saturation value and remains significantly above it. Combining Eq. (6) and Eq. (9), the minimum frame potential is attained at around
@@
Figure 6: Experimental results on frame potential versus the number of mid-circuit measurements. The dynamics are generated by $H_{\rm Ising}$ and executed experimentally on the QPU of "ibm_torino". Error bars denote the standard deviation across independent data ensembles. The total evolution time is $T=3/J_{zz}$. The system contains 3 system qubits coupled to 3 bath qubits. Other parameters are $J_{z}=0.9J_{zz}$ and $J_{x}=1.4J_{zz}$.
VI Experimental results
In experiments, mid-circuit measurement is crucial for implementing quantum error correction [google2025quantum] and realizing dynamical circuits [Baumer2024, Baumer20242]. Such capabilities enable scaling to larger system sizes and support more efficient quantum computation [cao2025].
We performed a proof-of-principle experiment on a small-scale system using the IBM Quantum Platform. Hamiltonian evolution is realized through a Suzuki–Trotter decomposition. The first-order frame potential $F^{(1)}$ can be extracted via full state tomography on the system qubits for the system size in our study (Appendix C). As shown in Fig. 6, the experimental data (red dots), exhibit good agreement with the theoretical prediction (blue solid line), with a finite reduction attributable to experimental noise, consistent with the corresponding noisy simulation (yellow dash-dotted line).
VII Conclusion
We propose a protocol based on multiple Hamiltonian evolutions, mid-circuit measurements, and resets to realize holographic deep thermalization. This approach reduces the number of required bath (ancilla) qubits and relaxes the need for large-scale qubit entanglement in experiments aimed at generating quantum randomness.
@@ Acknowledgements.
The authors acknowledge discussions with Bingzhi Zhang. The project is supported by Office of Naval Research Grant No. N00014-23-1-2296 and DARPA HR0011-24-9-0362. QZ also acknowledges support from NSF (OMA-2326746, 2350153, CCF-2240641), AFOSR MURI FA9550-24-1-0349, DARPA HR0011-24-9-0362, Halliburton Company, an unrestricted gift from Google.
@@ Appendix B Zeno effect
In the following, we provide the derivation of Eq. (9) from the main text. We consider Hamiltonian evolution in a composite system where a subset of qubits is designated as a bath and subject to repeated measurement and reset. In general, the Hamiltonian can be decomposed into three parts
@@
$=$ $H_{{\rm s}0}+H_{{\rm b}}+H_{{\rm c}0}\penalty 10000\ ,$
where $H_{{\rm s}},H_{{\rm b}}$ and $H_{{\rm c}}$ correspond to the system, bath and interaction Hamiltonian. In the second step, we introduced a modified system and interaction Hamiltonian
@@
$H_{{\rm c}0}=H_{\rm c}-\langle\phi_{0}|H_{\rm c}|\phi_{0}\rangle\penalty 10000\ ,$ (16)
@@ B.1 Low order Zeno effect
In the general case, the power is $\alpha=1$. The measurement basis is changed under the operation $H_{{\rm c}0}$, i.e., $\left[H_{{\rm c}0},|\phi_{m}\rangle\langle\phi_{m}|\right]\neq 0$. We can decompose the system wave function in the orthogonal basis
@@
The probability of the bath remaining in the state $|\phi_{0}\rangle$ is
$P_{0}=1-\sum_{m\neq 0}P_{m}\approx 1-\Delta t^{2}\sum_{m\neq 0,l^{\prime}}\overline{|\omega_{l^{\prime},m}|^{2}}.$ (20)
@@
$>$ ${\rm Exp}\left(-\frac{2T^{2}}{n}\sum_{m\neq 0,l^{\prime}}\overline{|\omega_{l^{\prime},m}|^{2}}\right).$
In this case, the coefficient is $c_{H}=2\sum_{m\neq 0,l^{\prime}}\overline{|\omega_{l^{\prime},m}|^{2}}$.
B.2 High-order Zeno effect
In the special case, the power is $\alpha=3$ when the measurement basis $|\phi_{m}\rangle$} is invariant under the first order action of $H_{\rm c0}$, i.e., $\left[H_{\rm c0},|\phi_{m}\rangle\langle\phi_{m}|\right]=0$. Under this symmetry condition, we obtain a tighter lower bound on the frame potential.
For convenience in the derivation, the bath Hamiltonian is decomposed into a component diagonal in the measurement basis and an off-diagonal term,
$H_{{\rm b}}=H_{{\rm bm}}+V\penalty 10000\ ,$ (22)
@@ B.2 High-order Zeno effect
In the limit of short time evolution $\Delta t\rightarrow 0$, the unitary could be written as [WangYan2024]
@@
$|\Psi_{t+\Delta t}\rangle$ $\approx$ $e^{-iH_{{\rm c0}}\frac{\Delta t}{2}}e^{-iH_{{\rm s0}}\Delta t-iH_{{\rm b}}\Delta t}e^{-iH_{{\rm c0}}\frac{\Delta t}{2}}|\psi_{t}\rangle|\phi_{0}\rangle$ (24)
@@
$|\Psi_{t+\Delta t}\rangle\approx\sum_{m}e^{-i\langle\phi_{m}|H_{\rm c0}|\phi_{m}\rangle\frac{\Delta t}{2}}c_{m}|\psi_{t+\Delta t}\rangle|\phi_{m}\rangle.$ (25)
In the first-order time-dependent perturbation theory, we consider only the states directly coupled to the initial state. The corresponding amplitude for $m\neq 0$ is given by
@@
$P_{m}$ $\approx$ $4\left|\frac{\langle\phi_{m}|V|\phi_{0}\rangle}{E_{0}-E_{m}}\right|^{2}\sin^{2}\frac{\left(E_{0}-E_{m}\right)\Delta t}{2}\penalty 10000\ ,$ (26)
and $P_{0}=1-\sum_{m\neq 0}P_{m}$.
@@
where $|\psi_{T}\rangle$ is the system wave function when the measurement results remain $|\phi_{0}\rangle$ throughout the entire process.
@@
We define the time scale $\mu_{m}$ as half the difference between the largest and smallest eigenvalues of $\langle\phi_{m}|H_{\rm c0}|\phi_{m}\rangle$. The frame potential then simplifies to
$F^{(K)}$ $>$ $\left(P_{0}+\sum_{m\neq 0}P_{m}\cos^{2K}\mu_{m}\Delta t\right)^{2n}$ (28)
$>$ $\left(1-\sum_{m}\left|\langle\phi_{m}|V|\phi_{0}\rangle\right|^{2}K\mu_{m}^{2}\Delta t^{4}\right)^{2n}$
$>$ ${\rm Exp}\left(-\frac{2KT^{4}}{n^{3}}\sum_{m}\left|\langle\phi_{m}|V|\phi_{0}\rangle\right|^{2}\mu_{m}^{2}\right)\penalty 10000\ .$
Because $\mu_{m}=0$ whenever $\langle\phi_{m}|H_{\rm c0}|\phi_{m}\rangle=0$, only a small subset of intermediate states contributes. For the Ising Hamiltonian $H_{\rm Ising}$, this expression reduces to $c_{H}=2J_{x}^{2}J_{zz}^{2}$.
Appendix C Experiment Methods
Figure 8: Experimental quantum circuit diagram implemented on the IBM Quantum Platform for $n=3$.
We implement our protocol on the IBM Quantum Platform, executing it on the “ibm_torino” QPU. The experimental circuit, shown in Fig. 8, uses $n_{\rm all}=6$ qubits and one classical bit. Among them, $n_{\rm s}=3$ qubits form the system and evolve coherently throughout the process which are tomographed at the end. The remaining $n_{\rm b}=3$ qubits serve as the bath and are measured and reset at every iteration. The classical bit is utilized to enable the reset of the bath after measurements via the following procedure. Each measurement outcome is written to the classical bit: if the outcome is 1, an $X$ gate is applied; otherwise the qubit is left unchanged.
Hamiltonian evolution is implemented using a symmetric Suzuki–Trotter decomposition, which reduces sensitivity to Trotter ordering. In our experiment, the number of Trotter repetitions is 5 for $n=1$, 2 for $n=2$ and 1 for larger $n$ in our experiment.
For small system sizes, we perform full Pauli tomography. The reduced density matrix is expanded as
$\rho_{\rm s}=\sum_{j}p_{j}\sigma_{j}\penalty 10000\ ,$ (29)
where each coefficient $p_{j}$ is obtained by measuring in the corresponding Pauli basis. The first-order frame potential then follows as
$F^{(1)}={\rm Tr}(\rho_{\rm s}^{2})=N_{\rm s}\sum_{j}p_{j}^{2}\penalty 10000\ .$ (30)
For the larger system size, we use classical-shadow tomography. A random local unitary $U_{j}$ is applied before measuring in the computational basis, yielding a bitstring $|b_{j}\rangle$. The corresponding snapshot is
$\rho_{j}=(N_{\rm s}+1)U_{j}^{\dagger}|b_{j}\rangle\langle b_{j}|U_{j}-I\penalty 10000\ ,$ (31)
which can be computed efficiently. The frame potential is estimated using the unbiased shadow estimator
$F^{(1)}={\rm Tr}(\rho_{\rm s}^{2})\approx\frac{\sum_{j<k}{\rm Tr}(\rho_{j}\rho_{k})}{N_{\rm samp}(N_{\rm samp}-1)}\penalty 10000\ ,$ (32)
where $N_{\rm samp}$ is the number of snapshots. The snapshot budget is small, satisfying $N_{\rm samp}\ll N_{\rm s}$.
Higher-order frame potentials can, in principle, be accessed by running two identical circuits in parallel and performing a SWAP test to evaluate the overlap between their output states. However, this approach requires an experimental error rate roughly half that of the single-copy protocol.
The noisy simulations presented in Fig. 6 of the manuscript employ the built-in noise model of the IBM Quantum Platform, with all parameters directly extracted from the “ibm_torino” device. A subset of these parameters is listed in Tab. 1. The simulated architecture matches the experimental configuration.
Table 1: Selected device parameters of the QPU “ibm_torino”.
Parameter Label Median Value
[…]
Gate length $t_{\rm g}$ 68 ns
[…]
In the noisy intermediate-scale quantum (NISQ) regime, computational resources are severely constrained. Below we present an order-of-magnitude analysis adapted to contemporary superconducting architectures.
The total accumulated error rate $\varepsilon_{\rm a}$ can be obtained from individual errors $\varepsilon_{j}$, via $1-\varepsilon_{\rm a}=\prod_{j}\left(1-\varepsilon_{j}\right)$. Approximately, we have
$\ln\left(1-\varepsilon_{\rm a}\right)$ $=$ $\sum_{j}\ln\left(1-\varepsilon_{j}\right)\approx-\sum_{j}\varepsilon_{j}\penalty 10000\ ,$ (33)
where the index $j$ runs over all relevant noise channels. This linearized form provides a convenient approximation for estimating error budgets in the NISQ regime.
For highly entangled states for $n_{\rm s}$ qubits, the effective coherence time is reduced to
$T_{\rm c}\approx\frac{T_{\rm d}}{n_{\rm s}}\penalty 10000\ .$ (34)
The total evolution time required for a Trotterized implementation scales as
$T_{\rm Tro}\sim t_{\rm g}n_{\rm all}^{\zeta}n_{\rm Tro}\penalty 10000\ ,$ (35)
where $\zeta\in[0,1]$ quantifies the parallelizability of the circuit ($\zeta=0$ for fully sequential execution; $\zeta=1$ for fully parallel execution). The Trotter number is typically constrained by digitalization errors and scales as $n_{\rm Tro}\sim n_{\rm all}J_{zz}T$ [Childs2021, Ostmeyer_2023, Myers2023], where $n_{\rm all}$ is the total number of qubits. The total wall-clock duration of the protocol is then
$T_{\rm all}$ $=$ $T_{\rm Tro}+(n-1)n_{\rm b}^{\zeta}t_{\rm r}+n_{\rm s}^{\zeta}t_{\rm r}\penalty 10000\ .$ (36)
Gate errors accumulate approximately as
$\varepsilon_{\rm g}\sim\varepsilon_{2}n_{\rm all}n_{\rm Tro}\penalty 10000\ ,$ (37)
where single-qubit gate errors are neglected.
The measurement and reset error of the bath scales as $(n-1)n_{\rm b}$, while the readout error of the system is $n_{\rm s}\varepsilon_{\rm m}$.
Combining all contributions, the total error budget satisfies
$\ln\left(1-\varepsilon_{\rm a}\right)$ $\approx$ $-\frac{T_{\rm all}}{T_{\rm c}}-\varepsilon_{\rm g}-\left((n-1)n_{\rm b}+n_{\rm s}\right)\varepsilon_{\rm m}.$ (38)
The first term originates from the finite coherence time, the second from accumulated gate errors, and the third from measurement errors. In the highly parallel operation $\zeta=0$ and small number of mid-circuit measurement $n\rightarrow 1$, the total error is dominated by the accumulated gate errors.
For first-order Suzuki–Trotter evolution under $H_{\rm Ising}$, the accumulated error satisfies
$\varepsilon_{\rm a}\gtrsim 2\varepsilon_{2}n_{\rm all}^{2}J_{zz}T$ (39)
or equivalently
$J_{zz}T\lesssim\frac{\varepsilon_{\rm a}}{2n_{\rm all}^{2}\varepsilon_{2}}.$ (40)
Substituting the experimental parameters yields the constraint $J_{zz}T\lesssim 5$. Combining this with Eq. (12) of the manuscript then implies that the optimal number of mid-circuit measurements is bounded by $n_{\gamma}\lesssim 4$.
Future improvements in random-state generation will require reduced gate error rates, shorter gate durations, and longer qubit coherence times. Most importantly, native Hamiltonian evolution and parallelized gate execution at the hardware level would significantly loosen these constraints.